OpenStudy (anonymous):

y=ln l4x^2-9xl Find the derivative of function

OpenStudy (anonymous):

we will use chain rule over here \[\frac{ dy }{ dx } = \frac{ 1 }{ 4x^2-9x } *\frac{ d }{ dx }(4x^2-9x)\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 4x^2-9x }* (8x-9)\]

OpenStudy (anonymous):

Whats the next step And how did you got 8x?

OpenStudy (anonymous):

ok look at this step\[4\frac{ d }{ dx }x^2 = 4*2x = 8x\] is this step clear to you?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do u know \[\frac{ d }{ dx } x^n = n*x^{n-1}\] ????????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so i used this concept in differentiating 4x^2

OpenStudy (anonymous):

okay